Davis Inc., operates a downhill ski area near Champaign, Illinois. An all-day adult lift ticket can be purchased for $85. Adult customers can purchase a season pass that entitles the pass holder to ski any day during the season, which typically runs from December 1 through April 30. Davis Inc. expects its season pass holders to use their passes equally throughout the season. The company's fiscal year ends on December 31.

On November 6, 2016, Ima Skier purchased a season pass for $450.

What date(s) should Davis Inc. recognize revenue from the sale of its season passes?

November 6, 2016

November 6, 2016–December 31, 2016

December 1, 2016–April 30, 2017

April 30, 2017

2.Question 2

What is the amount of revenue that will be included in the Davis Inc. 2016 Income Statement related to the sale of the season pass to Ima Skier?

3.Question 3

What is the remaining balance that will be included in the Davis Inc. 2016 Balance Sheet related to the sale of the season pass to Ima Skier?

4.Question 4

Problem 2: Evaluating the Efficiency of Asset Management

Use the following information for questions 4-6.

The 2016 Income Statement of Illini Company reported net sales of $8 million, cost of goods sold of $4.8 million, and net income of $800,000. The following table shows the company's comparative Balance Sheets for 2016 and 2015:

2016 ($ in 000s)

2015 ($ in 000s)

Assets

Cash

$ 300

$ 380

Accounts receivable

700

500

Inventory

900

700

Property, plant, and equipment (net)

2,400

2,120

Total assets

$4,300

$3,700

Liabilities and Shareholders' Equity

Current liabilities

$ 960

$ 830

Bonds payable

1,200

1,200

Paid-in capital

1,000

1,000

Retained earnings

1,140

670

Total liabilities and shareholders' equity

$4,300

$3,700

The industry averages for Illini’s line of business are:

  • Inventory turnover: 5 times
  • Average collection period: 25 days
  • Asset turnover: 1.8 times

In the following questions, assess Illini's asset management relative to its industry by calculating the key three activity ratios above.

What is the inventory turnover ratio for Illini Company for 2016?

5.Question 5

What is the asset turnover ratio for Illini Company for 2016?

6.Question 6

What is the average collection period for Illini Company’s accounts receivable? (Round to the nearest tenth days)
Answer #1

Answer 1)

Davis Inc. should recognize revenue from the sale of its season passes from December 1, 2016, to April 30, 2017, because the season pass is for the period starting from December 1 to April 30.

Answer 2)

The amount of revenue that will be included in the Davis Inc. 2016 Income Statement related to the sale of the season pass to Ima Skier is calculated as under:-

Davis Inc fiscal year-end on 31 December 2016, Pass is from Dec 1 to April 30. So

Amount of revenue:- $450*1/5 = $90

Answer 3)

Amount that will be included in the Davis Inc. 2016 Balance Sheet related to the sale of the season pass to Ima Skier is calculated as under:-

$450*4/5= $360

Answer 4)

Inventory turnover Ratio:- Cost of Goods Sold/Average Inventory = $4800000/$800000= 6

;- COGS:-$4.8 million

:- Average Inventory= (Opening nventory+Closing Inventory)/2

( $700000+$900000)/2= $800000

Answer 5) Asset turnover Ratio:- Net sales/Average Assets= $4800000/$4000000= 1.2

Average Assets= (Opening Assets + Closing Assets)/2

($3700000+$4300000)/2= $4000000

Answer 6) Average Collection Period:- Average Account receivable/Net credit sales* 360

= $600000/$8000000*360= 27 Days

:-     Average Account receivable= (Opening Acc Receivable+ Closing Acc Receivable)/2

($500000+$700000)/2= $600000
Answer #2
6. 27÷10=2.7
source: Mba
answered by: Neha
Answer #3


Answer 5) Asset turnover Ratio : Net sales/Average Assets = $8,000,000/$4,000,000 = 2

source: My brains
answered by: Ms. Imtiyaz
Answer #4

Davis Inc fiscal year-end on 31 December 2016, Pass is from Dec 1 to April 30. So

Amount of revenue:- $450*1/5 = $90


source: 360
answered by: magdalenemattia95@gmail.com